OpenStudy (jazilove):

PLEASE HELP, WILL MEDAL 7/13 divided by 5/39

satellite73 (satellite73):

same as \[\frac{7}{13}\times \frac{36}{5}\]

satellite73 (satellite73):

cancel first, multiply last

OpenStudy (jazilove):

A. 4 1/5 B. 5/21 C. 11 2/3 D. 21/5

OpenStudy (math&ing001):

Little typo there \[\frac{ 7 }{ 13 } * \frac{ 39 }{ 5 } \] Please put on a little effort and actually multiply and simplify to get your answer..

OpenStudy (jazilove):

wait so I multiply?

satellite73 (satellite73):

cancel first, multiply last

OpenStudy (math&ing001):

^